# Algorithmic Collusion And The Minimum Price Markov Game

### v2

@@ Abstract
To replicate the study, follow GitHub.
@@ 1 Introduction
In this paper, we introduce the Minimum Price Markov Game (MPMG), a model designed to provide a framework for studying emergent behavior in the context of minimum price auctions, which are commonly used in real-world public procurement. We draw on Robert Axelrod’s concept of “the emergence of cooperation among egoist” [22], exploring tacit coordination among isolated agents in the MPMG. Departing from the traditional game-theoretic Bayesian framework of first-price auctions, we consider a game of complete information based on the assumption of data availability. Indeed, public auction data typically includes the features of auctions and agents from past to ongoing contracts, as well as the associated bids [23]. This allows us to base the MPMG on the Prisoner’s Dilemma (PD), leveraging a structure that captures the essence of public procurement as a social dilemma—competitive by law but cooperative by nature. Indeed, the PD is well known for encapsulating the paradox of rationality central to our discussion on the study of emergent coordination among isolated agents in public markets. Furthermore, we test the MPMG using various state-of-the-art Multi-Agent Reinforcement Learning (MARL) methods and thereby show that the minimum price rule, in our setting, is generally difficult to achieve but far from impervious to tacit coordination among isolated decision-makers.
@@ 2 Related Work
The interplay among MDPs, auction design, and multi-agent learning has gained significant attention in recent research. Although the literature addressing algorithmic collusion within this framework is relatively sparse, our study draws inspiration from the growing body of work on integrating Stackelberg game models within multi-agent learning frameworks, MDPs for auction designs, and tacit coordination in social dilemmas. By reviewing these interconnected areas, we aim to highlight the advancements in understanding strategic interactions, optimal bidding strategies, and tacit coordination among agents. This overview highlights the pivotal role of MDPs and related methodologies in shaping contemporary economic and strategic decision-making processes.
@@
Stackelberg models, in which players move sequentially rather than simultaneously and therefore were historically used to study hierarchical or sequential leadership, are being increasingly used in various applications within the Markov game and multiagent learning framework [32]. The Stackelberg game is a staple in the analysis of various economic and strategic situations where the ability to move first or commit to a decision before one’s competitors can confer a strategic advantage. This includes industries with clear market leaders who can influence the market environment [33]. For example, RL agents have been trained to learn the Stackelberg equilibrium in a Markov decision process representing economic mechanisms for assigning items to individualistic agents in sequential-pricing stages [34, 35, 36].
@@
In scenarios where direct communication is either limited or non-existent, learning algorithms play a pivotal role in facilitating implicit cooperation and understanding competitive behaviors. For instance, learning methods that incorporate the anticipation of other agents’ learning processes have shown promise in fostering cooperative strategies that are robust against exploitation [41]. This is particularly evident in the Iterated Prisoner’s Dilemma, where strategies developed through evolutionary algorithms have been demonstrated to outperform traditional approaches [42]. Indeed, the contradictory nature of these games makes their study with RL agents quite challenging, as demonstrated in [43]. Moreover, the development of frameworks that enforce cooperation and resist collusion highlights the potential for creating stable cooperative environments in public goods games and other multi-agent settings [44, 45].
@@ 3 The Minimum Price Markov Game
Furthermore, in public procurement, there is a tension between the pursuit of personal gain and collective welfare. Firms strive to maximize their individual profits, while the state aims to allocate public funds efficiently and maintain market competition and fairness. From the agents’ perspective, individually rational choices made without cooperation can lead to suboptimal outcomes for all participants, as maximal profits are only achievable through cooperative behavior. This conflict between individual and collective rationality is the core characteristic of social dilemmas. The minimum price rule, with its mathematical elegance, does not diminish this inherent conflict; rather, it preserves this tension fully, maintaining the balance between self-interest and group benefit.
@@
In addition, the single-stage formulation must capture the core elements of market dynamics, which inherently unfold through repeated interactions. These characteristics must be maintained in its dynamic extension, such as in the iterated MPG and the Minimum Price Markov Game (MPMG).
In this section, we provide a formulation for the MPMG, which we believe is a reasonable approximation of the typical real-world minimum price-ruled public procurement process. Our formulation is grounded in assumptions that capture the core essence of this process, creating a conducive environment to observe how artificial learners dynamically shape their strategies in such dilemma. The MPMG is founded on a static normal-form game, the Minimum Price Game (MPG), which we will define first.
@@ 3.1 The Minimum Price Game: A Single-Stage Formulation
Let $b_{weak}$ and $b_{strong}$ be the bids in an auction within a heterogeneous duopoly. Naturally, and according to (1), we have $b_{strong}<b_{weak}$. Assuming the payoffs are determined by a minimum price rule, i.e., $u_{i}=b_{i}$ if $b_{i}=\min(b_{weak},b_{strong})$ and $u_{i}=0$ otherwise, we obtain the following bimatrix game:
@@
$u_{i}(\textit{FP},k>0)>u_{i}(\textit{CP},k>0).$
In fact, we can set $u_{i}(\textit{CP},k>0)=0$, meaning that when some players defect (FP) while others cooperate (CP), the cooperating agents receive no payoff, while defecting agents earn a share based on their market strengths relative to the defecting opponents. Let $\Omega\subseteq\mathcal{N}$ denote the set of agents playing FP, and let the total market power of the defecting agents be $\beta_{\Omega}=\sum_{j\in\Omega}\beta_{j}$. Assuming symmetric play where $\beta_{\Omega}=1$, the payoffs in the general heterogeneous case are given in Table 1.
@@
Strategy Profile $\mathbf{k=0}$ $\mathbf{k>0}$ $\mathbf{k=n-1}$
@@
The pairwise relation in terms of incentive factors between agent $i$ and agent $o$ is given by
$\gamma_{i}=\frac{\beta_{o}}{\beta_{i}}\gamma_{o}.$ (3)
@@
This sharing mechanism allows for heterogeneous agents to coexist while preserving the dynamic stemming from the minimum price. In fact, the power parameter $\beta$ not only represents the ability to cut costs but also the work capacity. As strong agents are more rewarded than weak ones in symmetrical plays (see Table 1), the payoff structure of the MPG emulates the natural turnover in auction participation of heterogeneous agents observed in markets, making the coexistence between weak and strong agents possible. Indeed, a higher payoff mathematically represents the higher frequency of wins by stronger agents in a strict minimum price-ruled process. When the setting simplifies to a homogeneous game, the sharing mechanism in symmetrical plays smooths out a random turnover between agents of equal power, as shown in Example 3.
@@ 3.2 Markov Game Formulation
In contrast to iterated games, Markov games evolve over multiple stages, with states dependent on previous actions. This dynamic nature better captures the complexities of real-world markets where strategies evolve over time, especially in the context of behavioral emergence, which may heavily rely on self-strengthening trends. The MPMG extends the framework of the MPG by providing the advantage of the Markov property, which relies on a state space that encodes information about opponents’ past plays and characteristics.
@@
Markov game. A Markov game, also known as a stochastic game, is a generalization of a Markov Decision Process (MDP) to multiple interacting agents contained in the set $N$. Each player $i$ has an associated action space $\mathcal{A}_{i}$. The game is defined over a common state space $\mathcal{S}$. The dynamics of the game are captured by a transition function $T:\left(\bigtimes_{i\in N}\mathcal{A}_{i}\right)\times\mathcal{S}\rightarrow[0,1]$, which determines the probability of moving from one state to another based on the actions of all players. Each player $i$ has a reward function $R_{i}:\left(\bigtimes_{i\in N}\mathcal{A}_{i}\right)\times\mathcal{S}\rightarrow\mathbb{R}$, defining the payoff $r_{i}$ for player $i$ based on the current state and actions taken by all players. The objective for each player $i$ is to maximize the expected sum of discounted rewards, expressed as $E\left[\sum_{k=0}^{\infty}\gamma^{k}r_{i,t+k}\right]$, where $\gamma$ is the discount factor that values present rewards over future rewards. A key feature of an MDP, and therefore of a Markov game, is the Markov property, which asserts that future states depend only on the current state and the actions taken by the agents, not on the sequence of events that preceded it, ensuring that past states are irrelevant once the current state is known.
@@
The choice of state space in a Markov game is often at the discretion of the implementer, depending on the specifics of the problem and the goals of the implementation. However, for our purposes, we define the MPMG with core variables that capture essential aspects of the game. Specifically, these sets include all players’ information on action frequencies $\mathbf{\bar{a}}^{t}=\{\bar{a}_{j}^{t}\}_{j\in\mathcal{N}}$, joint action frequencies $\mathbf{\bar{f}}^{t}=\{\bar{f}_{j}^{t}\}_{j\in\mathcal{N}}$, average rewards $\mathbf{\bar{r}}^{t}=\{\bar{r}_{j}^{t}\}_{j\in\mathcal{N}}$, and power parameters $\mathbf{\beta}=\{\beta_{j}\}_{j\in\mathcal{N}}$. Therefore, the state space at any time $t$ is $\mathcal{S}_{t}=\{\mathbf{\bar{a}}^{t},\mathbf{\bar{f}}^{t},\mathbf{\bar{r}}^{t},\mathbf{\beta}\}$, and is of size $2n+|\mathcal{A}|^{n}$. This minimal configuration for the state space aligns with Assumption 4. Furthermore, the power parameters also allow for the possibility of modeling a dynamic market in which agents could lose or gain market power.
@@
The payoff structure defined for the MPG translates to a reward function in the MPMG, emphasizing the inherent connection between the Markov game and its RL implementation. The reward function, together with the transition function, which defines how the game state evolves in response to actions, fundamentally shapes the dynamics and outcomes of the game. In fact, in the classical RL setting, the normalized reward function for agent $i$ at each iteration $t$ is expressed as
@@
We do not make this assumption formal, as this normalization does not affect the generality of the model but only simplifies the calculations. The immediate rewards for agent $i$ at nay time $t$ are given in table 1.
@@
The MPMG unfolds as an episodic game of arbitrary length (number of episodes), with each episode consisting of a single-step auction (Assumptions 1 and 2). Agents compete against each other in an environment representing the same entity in each episode (Assumption 5). Due to the Markov property, memory is encapsulated in the observation space $\mathcal{S}$ at any episode $t$. Agents use this observation space to learn their strategies, conditional on occurrences in $\mathcal{S}$, effectively playing a Markov game where the reward function, $R(a_{1},\dots,a_{n})$, is deterministic and based on the joint action. The transition probability function is given by $P(s^{\prime}|s,a)$, where $s^{\prime}$ is the new state and $s\in\mathcal{S}$. When the state space is not used for decision-making, the agents are effectively playing an iterated MPG.
@@ 4 Computer Experiments
Cooperation in multi-agent reinforcement learning (MARL) settings is well studied. However, this algorithmic cooperation mostly relies on explicit designs—such as shared joint action spaces [48], algorithmic communication [49], reputation-based models [50]—which are not suitable for our context, as these approaches emulate scenarios where agents deliberately bias their actions towards cooperation.
@@
We conducted a total of twenty-four experiments, combining six agent types with four MPMG configurations (2-player homogeneous, 5-player homogeneous, 2-player heterogeneous, and 5-player heterogeneous MPMG). As discussed in Section 3, $0\leq\sigma(\beta)\leq 1$ represents the standard deviation of the power parameters ($\beta$) distribution, with $\sigma(\beta)=0$ representing the homogeneous case. For experiments in the heterogeneous setting, we set $\sigma(\beta)=0.5$. Additionally, we set $\alpha=1.3$, meaning that the collusive bids were 30% higher than the fair bids. Each agent was trained on 100 auctions over 100 replications for statistical robustness.
@@ 4.1 Agent Representation
Multiagent Reinforcement Learning (MARL). MARL extends single-agent reinforcement learning (RL) techniques to settings with multiple agents learning concurrently in a shared environment. In RL, an agent interacts with its environment to maximize cumulative rewards, where key components include agent, environment, state, action, reward, policy, and value function. The agent, acting as the decision-maker, responds to the environment, which consists of external factors and situations represented as states. Actions are possible decisions the agent can take, evaluated by rewards as feedback on their success. The policy is the agent’s strategy for selecting actions based on the current state, while the value function estimates future rewards to determine the desirability of states or actions. In MARL, multiple agents simultaneously learn and update their policies, causing the environment to appear non-stationary from any single agent’s perspective. This dynamic leads to more complex interactions compared to single-agent RL, requiring robust techniques to address the challenges of Markov games and multiagent systems in general.
@@
Methods based on Q-learning [55] maintain a state-action pair values, named Q-values, based on the state-action pair expected reward. Decision making is then conducted following the highest Q-values associated to state-action pair in question. In our experiments, we used the Double Dueling Deep Q-network (D3QN) algorithm [56]. In the D3QN, a multi-layer perceptron is used to approximate the Q-values, and the Double Q-learning approach is employed to reduce overestimation bias by maintaining two sets of Q-value estimates. The Dueling architecture, on the other hand, decomposes the Q-value into two separate streams representing the state-value function and the advantage function, which helps the model learn better representations of the value of being in a given state, independent of the specific actions.
@@
Unlike Q-learning, policy gradient methods focus on optimizing a parameterized policy $\pi_{\theta}(a|s)$ directly by gradient ascent on expected rewards [55]. In the actor-critic framework, two function approximators are employed. The actor, which is responsible for selecting actions based on the current policy, and the critic, which evaluates the action-value function to provide feedback on the quality of the chosen actions. The actor updates the policy in the direction suggested by the critic, while the critic learns to estimate the expected return, enabling a stable and efficient training process by reducing the variance of the policy gradient estimates. Here, we used Multi-Agent Proximal Policy Optimization (MAPPO) agents [57], that extend the Proximal Policy Optimization (PPO), a popular policy gradient method, to multi-agent settings. The core idea of PPO is to strike a balance between policy improvement and preventing drastic updates to the policy, which could harm performance. To achieve this, PPO uses a modified objective function that constrains how much the new policy is allowed to diverge from the old policy during an update.

### v3

@@ Abstract
To replicate the study, follow replication repo (GitHub).
MPMG is implemented in a Python package; see PyPI or GitHub.
@@ 1 Introduction
In this paper, we introduce the Minimum Price Markov Game (MPMG), a model designed to provide a framework for studying emergent behavior in the context of minimum price auctions, commonly used in real-world public procurement. We draw on Robert Axelrod’s concept of “the emergence of cooperation among egoist” [22], exploring tacit coordination among isolated agents in the MPMG. Departing from the traditional game-theoretic Bayesian framework of first-price auctions, we consider a game of complete information based on the assumption of data availability. Indeed, public auction data typically includes the features of auctions and agents from past to ongoing contracts, as well as the associated bids [23]. This allows us to base the MPMG on the Prisoner’s Dilemma (PD), leveraging a structure that captures the essence of public procurement as a social dilemma—competitive by law but cooperative by nature. Indeed, the PD is well known for encapsulating the paradox of rationality central to our discussion on the study of emergent coordination among isolated agents in public markets. Furthermore, we test the MPMG using various state-of-the-art Multi-Agent Reinforcement Learning (MARL) methods and thereby show that the minimum price rule, in our setting, is generally robust but not impervious to tacit coordination among isolated decision-makers.
@@ 2 Related Work
The interplay among Markov Decision Process (MDP), auction design, and multi-agent learning has gained significant attention in recent research. Although the literature addressing algorithmic collusion within this framework is relatively sparse, our study draws inspiration from the growing body of work on integrating Stackelberg game models within multi-agent learning frameworks, MDP for auction designs, and tacit coordination in social dilemmas. By reviewing these interconnected areas, we aim to highlight the advancements in understanding strategic interactions, optimal bidding strategies, and tacit coordination among agents. This overview highlights the pivotal role of MDP and related methodologies in shaping contemporary economic and strategic decision-making processes.
@@
Stackelberg models, in which players move sequentially rather than simultaneously and therefore were historically used to study hierarchical or sequential leadership, are being increasingly used in various applications within the Markov game and multiagent learning framework [32]. The Stackelberg game is a staple in the analysis of various economic and strategic situations where the ability to move first or commit to a decision before one’s competitors can confer a strategic advantage. This includes industries with clear market leaders who can influence the market environment [33]. For example, Reinforcement Learning (RL) agents have been trained to learn the Stackelberg equilibrium in a Markov decision process representing economic mechanisms for assigning items to individualistic agents in sequential-pricing stages [34, 35, 36].
@@
In scenarios where direct communication is either limited or non-existent, learning algorithms play a pivotal role in facilitating implicit cooperation and understanding competitive behaviors. For instance, learning methods that incorporate the anticipation of other agents’ learning processes have shown promise in fostering cooperative strategies that are robust against exploitation [41]. This is particularly evident in the Iterated Prisoner’s Dilemma, where strategies developed through evolutionary algorithms have shown to outperform traditional approaches [42]. Indeed, the contradictory nature of these games makes their study with RL agents quite challenging, as demonstrated in [43]. Moreover, the development of frameworks that enforce cooperation and resist collusion highlights the potential for creating stable cooperative environments in public goods games and other multi-agent settings [44, 45].
@@ 3 The Minimum Price Markov Game
Furthermore, in public procurement, there is a tension between the pursuit of personal gain and collective welfare. Firms strive to maximize their individual profits, while the state aims to allocate public funds efficiently and maintain market competition and fairness. From the firms’ perspective, individually rational choices made without cooperation can lead to suboptimal outcomes for all participants, as maximal profits are only achievable through cooperative behavior. This conflict between individual and collective rationality is the core characteristic of social dilemmas. The minimum price rule, with its mathematical elegance, does not diminish this inherent conflict; rather, it preserves this tension fully, maintaining the balance between self-interest and group benefit.
@@
In addition, the single-stage formulation must capture the core elements of market dynamics, which inherently unfold through repeated interactions. These characteristics must be maintained in its dynamic extension, such as in the iterated Minimum Price Game (MPG) and the Minimum Price Markov Game (MPMG).
In this section, we provide a formulation for the MPMG, that we believe is a reasonable approximation of the typical real-world minimum price-ruled public procurement process. Our formulation is rooted in assumptions that capture the core of this process, creating a conducive environment to observe how artificial learners dynamically shape their strategies in such dilemma. The MPMG is founded on a static normal-form game, the MPG, which we will define first.
@@ 3.1 The Minimum Price Game: A Single-Stage Formulation
Let $b_{weak}$ and $b_{strong}$ be the bids in an auction within a heterogeneous duopoly. Naturally, and according to (1), we have $b_{strong}<b_{weak}$. Assuming the payoffs $u$ are determined by a minimum price rule, i.e., $u_{i}=b_{i}$ if $b_{i}=\min(b_{weak},b_{strong})$ and $u_{i}=0$ otherwise, we obtain the following bimatrix game:
@@
$u_{i}(\textit{FP},k>0)>u_{i}(\textit{CP},k>0).$ (3)
In fact, because of the minimum price rule, we need to make sure that the cooperating agents receive payoffs of 00 when some other players defect. Also, we set a sharing mechanism where defecting agents earn a share based on their market strengths relative to the defecting opponents. Therefore, let $\Omega\subseteq\mathcal{N}$ denote the set of agents playing FP, and let the total market power of the defecting agents be $\beta_{\Omega}=\sum_{j\in\Omega}\beta_{j}$. This means that when agent $i$ plays FP, its payoff,
$\frac{\beta_{i}}{\beta_{\Omega}}\cdot b_{i},$
depends on $\beta_{\Omega}$. Note that in this case, $\beta_{\Omega}=1$ when $k=n-1$, and $\beta_{\Omega}=\beta_{i}$ when $k=0$. When agent $i$ plays CP, then its payoff is either 00 if $k>0$ or
$\alpha\cdot\beta_{i}\cdot b_{i},$
if $k=0$. Table 1 summarizes the individual payoffs for agent $i$ in all scenarios.
@@
Strategy Profile $\mathbf{k=0}$ $\mathbf{n-1>k>0}$ $\mathbf{k=n-1}$
@@
The payoff function presented in Table 1 respects (2) and (3) since,
$b_{i}<\min_{j\neq i}\alpha\cdot b_{j}\quad\forall i,$
and
$\frac{\beta_{i}}{\beta_{\Omega}}\cdot b_{i}>0.$
@@
The pairwise relation in terms of incentive factors between agent $i$ and an opponent $j$ is given by
$\gamma_{i}=\frac{\beta_{j}}{\beta_{i}}\gamma_{j}.$ (5)
@@
Our sharing mechanism, using $\beta_{\omega}$, allows for heterogeneous agents to coexist while preserving the dynamic stemming from the minimum price. In fact, the power parameter $\beta$ not only represents the ability to cut costs but also the work capacity. As strong agents are more rewarded than weak ones in symmetrical plays (see Table 1), the payoff structure of the MPG emulates the natural turnover in auction participation of heterogeneous agents observed in markets, making the coexistence between weak and strong agents possible. Indeed, a higher payoff mathematically represents the higher frequency of wins by stronger agents in a strict minimum price-ruled process. When the setting simplifies to a homogeneous game, the sharing mechanism in symmetrical plays smooths out a random turnover between agents of equal power, as shown in Example 3.
@@ 3.2 Markov Game Formulation
In contrast to iterated games, Markov games evolve over multiple stages, with states dependent on previous actions. This dynamic nature better captures the complexity of real-world markets where strategies evolve over time, especially in the context of behavioral emergence, which may heavily rely on self-strengthening trends. The MPMG extends the framework of the MPG by providing the advantage of the Markov property, that relies on a state space that encodes information about opponents’ past plays and characteristics.
@@
Markov game. A Markov game, also known as a stochastic game, is a generalization of a MDP to multiple interacting agents contained in the set $N$. Each player $i$ has an associated action space $\mathcal{A}_{i}$. The game is defined over a common state space $\mathcal{S}$. The dynamics of the game are captured by a transition function $T:\left(\bigtimes_{i\in N}\mathcal{A}_{i}\right)\times\mathcal{S}\rightarrow[0,1]$, which determines the probability of moving from one state to another based on the actions of all players. Each player $i$ has a reward function $R_{i}:\left(\bigtimes_{i\in N}\mathcal{A}_{i}\right)\times\mathcal{S}\rightarrow\mathbb{R}$, defining the payoff $r_{i}$ for player $i$ based on the current state and actions taken by all players. The objective for each player $i$ is to maximize the expected sum of discounted rewards, expressed as $E\left[\sum_{t=0}^{\infty}\gamma^{t}r_{i,t}\right]$, where $\gamma$ is the discount factor that values present rewards over future rewards. A key feature of an MDP, and therefore of a Markov game, is the Markov property, which asserts that future states depend only on the current state and the actions taken by the agents, not on the sequence of events that preceded it, ensuring that past states are irrelevant once the current state is known.
@@
The choice of state space in a Markov game is often at the discretion of the implementer, depending on the specifics of the problem and the goals of the implementation. However, for our purposes, we define the MPMG with core variables that capture essential aspects of the game. Specifically, these sets include all players’ information on action frequencies $\mathbf{\bar{a}}^{t}=\{\bar{a}_{i}^{t}\}_{i\in\mathcal{N}}$, joint action frequencies $\mathbf{\bar{f}}^{t}=\{\bar{f}_{i}^{t}\}_{i\in\mathcal{N}}$, average rewards $\mathbf{\bar{r}}^{t}=\{\bar{r}_{i}^{t}\}_{i\in\mathcal{N}}$, and power parameters $\mathbf{\beta}=\{\beta_{i}\}_{i\in\mathcal{N}}$. Therefore, the state space at any time $t$ is $\mathcal{S}_{t}=\{\mathbf{\bar{a}}^{t},\mathbf{\bar{f}}^{t},\mathbf{\bar{r}}^{t},\mathbf{\beta}\}$, and is of size $2n+|\mathcal{A}|^{n}$. This minimal configuration for the state space aligns with Assumption 4. Furthermore, the power parameters also allow for the possibility of modeling a dynamic market in which agents could lose or gain market power.
@@
The payoff structure defined for the MPG leads to a reward function in the MPMG, emphasizing the inherent connection between the Markov game and its RL implementation. The reward function, together with the transition function, which defines how the game state evolves in response to actions, fundamentally shapes the dynamics and outcomes of the game. In fact, in the classical RL setting, the normalized reward function for agent $i$ at each iteration $t$ is expressed as
@@
We do not make this assumption formal, as this normalization does not affect the generality of the model but only simplifies the calculations. The immediate rewards for agent $i$ at any time $t$ are given in Table 1.
@@
The MPMG unfolds as an episodic game of arbitrary length (number of episodes), with each episode consisting of a single-step auction (Assumptions 1 and 2). Agents compete against each other in an environment representing the same entity in each episode (Assumption 5). Due to the Markov property, memory is encapsulated in the observation space $\mathcal{S}$ at any episode $t$. Agents use this observation space to learn their strategies, effectively playing a Markov game where the reward function, $R(a_{1},\dots,a_{n})$, is deterministic and based on the joint action. The transition probability function is given by $P(s^{\prime}|s,a)$, where $s^{\prime}$ is the new state and $s\in\mathcal{S}$. When the state space is not used for decision-making, the agents are effectively playing an iterated MPG.
@@ 4 Computer Experiments
Cooperation in MARL settings is well studied. However, this algorithmic cooperation mostly relies on explicit designs—such as shared joint action spaces [48], algorithmic communication [49], reputation-based models [50]—which are not suitable for our context, as these approaches emulate scenarios where agents deliberately bias their actions towards cooperation.
@@
We conducted a total of twenty-four experiments, combining six agent types with four MPMG configurations (2-player homogeneous, 5-player homogeneous, 2-player heterogeneous, and 5-player heterogeneous MPMG). As discussed in Section 3, $0\leq\sigma(\beta)\leq 1$ represents the standard deviation of the power parameters ($\beta$) distribution, with $\sigma(\beta)=0$ representing the homogeneous case. For experiments in the heterogeneous setting, we set $\sigma(\beta)=0.5$. Additionally, we set $\alpha=1.3$, meaning that the collusive bids are 30% higher than the fair bids. Each agent was trained on 100 auctions over 100 replications for statistical robustness.
@@ 4.1 Agent Representation
Multiagent Reinforcement Learning. MARL extends single-agent RL techniques to settings with multiple agents learning concurrently in a shared environment. In RL, an agent interacts with its environment to maximize cumulative rewards, where key components include agent, environment, state, action, reward, policy, and value function. The agent, acting as the decision-maker, responds to the environment, which consists of external factors and situations represented as states. Actions are possible decisions the agent can take, evaluated by rewards as feedback on their success. The policy is the agent’s strategy for selecting actions based on the current state, while the value function estimates future rewards to determine the desirability of states or actions. In MARL, multiple agents simultaneously learn and update their policies, causing the environment to appear non-stationary from any single agent’s perspective. The dynamics lead to more complex interactions compared to single-agent RL, requiring robust techniques to address the challenges of Markov games and multiagent systems in general.
@@
Methods based on Q-learning [55] maintain state-action pair values, named Q-values, based on the state-action pair expected reward. Decision making is then conducted following the highest Q-values associated with the state-action pair in question. In our experiments, we used the Double Dueling Deep Q-network (D3QN) algorithm [56]. In the D3QN, a multi-layer perceptron is used to approximate the Q-values, and the Double Q-learning approach is employed to reduce overestimation bias by maintaining two sets of Q-value estimates. The Dueling architecture, on the other hand, decomposes the Q-value into two separate streams representing the state-value function and the advantage function, which helps the model learn better representations of the value of being in a given state, independent of the specific actions.
@@
Unlike Q-learning, policy gradient methods focus on optimizing a parameterized policy $\pi_{\theta}(a|s)$ directly by gradient ascent on expected rewards [55]. In the actor-critic framework, two function approximators are employed. The actor, which is responsible for selecting actions based on the current policy, and the critic, which evaluates the action-value function to provide feedback on the quality of the chosen actions. The actor updates the policy in the direction suggested by the critic, while the critic learns to estimate the expected return, enabling a stable and efficient training process by reducing the variance of the policy gradient estimates. Here, we used Multi-Agent Proximal Policy Optimization (MAPPO) agents [57] that extend the Proximal Policy Optimization (PPO), a popular policy gradient method, to multi-agent settings. The core idea of PPO is to strike a balance between policy improvement and preventing drastic updates to the policy, which could harm performance. To achieve this, PPO uses a modified objective function that constrains how much the new policy is allowed to diverge from the old policy during an update.